# MonoScale: Scaling Multi-Agent System with Monotonic Improvement

### v1

@@ 7 Conclusion
We study a common failure mode in open multi-agent systems under continual expansion: cold-start misrouting that leads to performance collapse. We propose MonoScale, which, upon each new agent integration, proactively collects execution-grounded evidence via agent-conditioned warm-up probes and distills both successes and failures into auditable, rollbackable natural-language routing memories. On the theory side, we formalize expansion as a contextual bandit problem and establish a stage-wise monotonic non-degradation guarantee under conservative fallback and trust-region constraints. Experiments on GAIA and Humanity’s Last Exam show that MonoScale delivers stable gains as the agent pool grows from 3 to 10, substantially mitigating the degradation of naive scale-up, and remains more robust in the presence of noisy agents. We hope this work provides a safe and controllable expansion protocol for large-scale, open agent onboarding in the future Agentic Web; we further aim to integrate MonoScale with retrieval-based routing to support million-scale agent ecosystems.
@@ A.3.4 Agent 10: Multimedia Agent (False Advertising)
Appendix B Proofs for Theoretical Analysis
B.1 Setup and Technical Assumptions
@@
Assumption B.1 (Non-interfering expansion).
@@
Assumption B.2 (Feasible conservative fallback).
@@
B.2 Proof of Lemma 4.1
@@
B.3 Proof of Lemma 4.2 (Expansion Bridge)
@@
where the second line uses that $\pi_{m_{k-1}}^{\uparrow}(y\mid x)=0$ for all $y\in\mathcal{Y}_{k}\setminus\mathcal{Y}_{k-1}$. By Assumption B.1, for all $y\in\mathcal{Y}_{k-1}$ we have $r_{k}(x,y)=r_{k-1}(x,y)$, hence
@@
B.4 Proof of Theorem 4.3 (Monotonicity Across Expansions)
Fix any $k\geq 1$. By Assumption B.2, the conservative fallback memory $m_{k-1}^{\uparrow}$ is included in the feasible edit set $\mathcal{U}_{k}(m_{k-1})$. Moreover, it trivially satisfies the KL constraint in (16) since
@@
By Assumption B.2 (realizability of the lifted policy), $\pi_{m_{k-1}^{\uparrow}}^{k}\equiv\pi_{m_{k-1}}^{\uparrow}.$ Thus (36) becomes
@@
Appendix C Additional Results and Analysis
C.1 GAIA Comprehensive Results
@@
MonoScale Scaling
Qwen-3-30B-A3B-Instruct(3 Agents) 49.06% 47.67% 26.92% 44.84%
5 Agents (w/ Memory) 52.83% 45.35% 42.31% 47.27% $\uparrow$+5.4%
7 Agents (w/ Memory) 60.38% 47.67% 30.77% 49.09% $\uparrow$+9.5%
@@
C.2 Malfunctioning Agent Pool
@@
Appendix D MonoScale/Expansion-Aware Memory Schema
D.1 Overview
@@ D.1.1 Key Common Items
D.2 Planning Phase Memory
D.2.1 JSON Template
@@
D.2.2 Planning-Specific Items
@@
D.3 Coordinating Phase Memory
D.3.1 JSON Template
@@
D.3.2 Coordinating-Specific Items
@@
Appendix E Case Study Trajectories: The Anatomy of Scaling Failures and Their Cure
@@
E.1 Case 1: The Trap of Cold-Start Misjudgment (Anagram Solving)
@@
E.2 Case 2: The Trap of Brittle Workflow Links (Bird Species ID)
@@
E.3 Case 3: The MonoScale Resolution (Scientific Discovery)

### v2

@@ 7 Conclusion
We study a common failure mode in open multi-agent systems under continual expansion: cold-start misrouting that leads to performance collapse. We propose MonoScale, which, upon each new agent integration, proactively collects execution-grounded evidence via agent-conditioned warm-up probes and distills both successes and failures into auditable, rollbackable natural-language routing memories. On the theory side, we formalize expansion as a contextual bandit problem and establish a stage-wise monotonic non-degradation guarantee under a non-interfering expansion assumption, conservative fallback, and trust-region constraints. Experiments on GAIA and Humanity’s Last Exam show that MonoScale delivers stable gains as the agent pool grows from 3 to 10, substantially mitigating the degradation of naive scale-up, and remains more robust in the presence of noisy agents. We hope this work provides a safe and controllable expansion protocol for large-scale, open agent onboarding in the future Agentic Web; we further aim to integrate MonoScale with retrieval-based routing to support million-scale agent ecosystems.
@@ A.3.4 Agent 10: Multimedia Agent (False Advertising)
Appendix B Expansion-Aware Memory Update Procedure
This section makes explicit the separation between the theoretical trust-region view used in Section 4 and the operational update procedure deployed in our system. We do not continuously optimize the router by directly solving the constrained surrogate objective; instead, with the router LLM kept frozen, each expansion stage performs a one-shot, conservative selection over a finite set of candidate memory states. Algorithm 1 summarizes the full procedure.
Algorithm 1 Expansion-Aware Memory Update at Stage $k$
0: deployed memory $m_{k-1}$; new agent $a_{k}$; agent cards for pool $\mathcal{S}_{k}$; frozen router LLM; rollouts per task $R$; KL radius $\delta$
0: updated memory $m_{k}$
1:Stage 1: Expansion-aware task synthesis
2: Synthesize warm-up tasks $\mathcal{T}_{k}^{\mathrm{syn}}$ conditioned on the agent card of $a_{k}$ via the planner–executor–validator loop
3:Stage 2: Evidence collection
4: Initialize experience buffer $\mathcal{B}_{k}\leftarrow\emptyset$
5:for each warm-up task $x\in\mathcal{T}_{k}^{\mathrm{syn}}$ do
6:  Sample $R$ orchestration rollouts $\{(y_{j},r_{j})\}_{j=1}^{R}$ under the current router $\pi^{k}_{m_{k-1}}$
7:  if $r_{j}=0$ for all $j$ then
8:   Discard $x$ {drop unsolvable or overly noisy tasks}
9:  else
10:   $\mathcal{B}_{k}\leftarrow\mathcal{B}_{k}\cup\{(x,y_{j},r_{j})\}_{j=1}^{R}$
11:  end if
12:end for
13:Stage 3: Candidate memory construction
14: Distill structured routing principles from the successful and failed traces in $\mathcal{B}_{k}$ using an LLM, forming candidate memories $\tilde{\mathcal{U}}_{k}$ (schema in Appendix E)
15: Reject candidates that semantically conflict with existing principles {semantic trust region}
16: Construct the conservative fallback memory $m_{k-1}^{\uparrow}$ that forbids invoking $a_{k}$
[…]
18:Stage 4: Conservative selection
[…]
20:if no candidate satisfies the trust-region constraint then
21:  $m_{k}\leftarrow m_{k-1}^{\uparrow}$ {conservative no-improvement step}
22:end if
23:return $m_{k}$
The trust-region/KL formulation enters only as the analytical tool that certifies stability in Theorem 4.3: candidate memories that would induce large behavioral shifts are either rejected by the semantic consistency check or rendered infeasible by the KL constraint, while the always-available conservative fallback $m_{k-1}^{\uparrow}$ guarantees that the selected update never underperforms the pre-expansion policy. In our implementation, we synthesize roughly $50$ warm-up tasks per newly added agent and sample $R=4$ rollouts per task; tasks that fail in all $R$ rollouts are discarded to limit the influence of unsolvable or overly noisy samples. The structured-memory schema used to represent each candidate state is detailed in Appendix E.
Appendix C Proofs for Theoretical Analysis
C.1 Setup and Technical Assumptions
@@
Assumption C.1 (Non-interfering expansion).
@@
Assumption C.2 (Feasible conservative fallback).
@@
C.2 Empirical Validation of the Technical Assumptions
Theorem 4.3 relies on two assumptions: non-interfering expansion (Assumption C.1) and a feasible conservative fallback (Assumption C.2). We empirically examine both.
Non-interfering expansion (Assumption C.1).
We check this assumption at two levels. (i) Decision stability. We fix the archived stage-(k1)(k\!-\!1) plans and compare the router’s selections before and after expansion; the selections overlap by more than $95\%$. Since we do not modify any existing agent’s configuration or tools during expansion, an agent’s expected reward on a given subtask is invariant across stages, so this high overlap already implies that the reward of the large majority of legacy plans is unchanged. (ii) Reward stability. Because routing overlap measures decision stability rather than reward directly, we additionally run a more direct test: we sample $100$ questions from HLE, fix the archived stage-(k1)(k\!-\!1) plans unchanged, and execute the same plans before and after expansion. The resulting accuracy is $19/100$ before expansion and $21/100$ after, a difference of only $2$ points that is not statistically significant. Together these results indicate that expansion does not systematically shift the reward of legacy plans, so Assumption C.1 holds approximately in our setting. The assumption is most likely to break under persistent environment drift—e.g., when the behavior or availability of external APIs changes during the expansion process—in which case legacy-plan rewards may shift for reasons orthogonal to agent expansion.
Feasible conservative fallback (Assumption C.2).
The fallback is realized by a natural-language memory entry that forbids invoking the newly added agent. Across our runs, an agent banned in this way is never selected by the router, confirming that the conservative fallback $m_{k-1}^{\uparrow}$ is not only theoretically available but also reliably realizable in practice.
C.3 Proof of Lemma 4.1
@@ C.2 Empirical Validation of the Technical Assumptions
C.4 Proof of Lemma 4.2 (Expansion Bridge)
@@
where the second line uses that $\pi_{m_{k-1}}^{\uparrow}(y\mid x)=0$ for all $y\in\mathcal{Y}_{k}\setminus\mathcal{Y}_{k-1}$. By Assumption C.1, for all $y\in\mathcal{Y}_{k-1}$ we have $r_{k}(x,y)=r_{k-1}(x,y)$, hence
@@
C.5 Proof of Theorem 4.3 (Monotonicity Across Expansions)
Fix any $k\geq 1$. By Assumption C.2, the conservative fallback memory $m_{k-1}^{\uparrow}$ is included in the feasible edit set $\mathcal{U}_{k}(m_{k-1})$. Moreover, it trivially satisfies the KL constraint in (16) since
@@
By Assumption C.2 (realizability of the lifted policy), $\pi_{m_{k-1}^{\uparrow}}^{k}\equiv\pi_{m_{k-1}}^{\uparrow}.$ Thus (36) becomes
@@
Appendix D Additional Results and Analysis
D.1 GAIA Comprehensive Results
@@
Naive Scale-up (Qwen-3-30B-A3B-Instruct, w/o Memory)
3 Agents 49.06% 47.67% 26.92% 44.84%
5 Agents 54.72% 43.02% 23.08% 43.64%
7 Agents 54.72% 40.70% 46.15% 46.06%
10 Agents 47.17% 44.19% 26.92% 42.42%
MonoScale (Qwen-3-30B-A3B-Instruct, w/ Memory)
3 Agents 49.06% 47.67% 26.92% 44.84%
5 Agents 52.83% 45.35% 42.31% 47.27% $\uparrow$+8.3%
7 Agents 60.38% 47.67% 30.77% 49.09% $\uparrow$+6.6%
@@
D.2 Ablation Study
To isolate the contribution of MonoScale’s key design choices, we ablate two components on GAIA with the Qwen3-30B-A3B router: (i) negative memory—the negative routing constraints distilled from failed warm-up traces—and (ii) customized warm-up—the agent-conditioned task synthesis. For the non-customized variant, we follow the cold-start setting of EvoRoute (Zhang et al., 2026a) and use $50$ standardized TaskCraft (Shi et al., 2025) tasks as generic warm-up data. As shown in Table 4, both components matter, especially at larger scales. Removing negative memory causes a sharp drop at $10$ agents ($55.2\%\rightarrow 40.0\%$), indicating that explicit “when not to route” constraints are essential for stable expansion. Replacing customized warm-up with generic tasks also lowers the final expanded accuracy ($55.2\%\rightarrow 48.5\%$). Overall, MonoScale’s gains stem from both negative routing memory and deployment-aware, agent-conditioned onboarding, rather than from additional warm-up or memory volume alone.
Table 4: Ablation study on GAIA (overall accuracy, Qwen3-30B-A3B router). w/o Negative Memory removes the negative routing constraints distilled from failed traces; w/o Customized Warm-up replaces agent-conditioned task synthesis with generic TaskCraft tasks. Both components contribute, most visibly at the largest agent pool.
Configuration 5 Agents 7 Agents 10 Agents
MonoScale (full) 47.3% 49.1% 55.2%
   w/o Negative Memory 45.5% 50.9% 40.0%
   w/o Customized Warm-up 50.9% 49.7% 48.5%
D.3 Cross-Router Memory Transfer
A natural question is whether the routing memory accumulated by MonoScale is tied to the specific router that produced it, or whether it captures router-agnostic knowledge about the agent pool. To probe this, we take the memory generated by the Qwen-3-30B-A3B-Instruct router during expansion and apply it directly, without any adaptation, to a different router backbone, Gemini-3-Flash, on GAIA across the same $3\rightarrow 10$ agent expansion. Table 5 compares this transferred setting (w/ Qwen memory) against Gemini-3-Flash running MonoScale with its own natively generated memory (w/ native memory).
Table 5: Cross-router memory transfer on GAIA (overall accuracy). w/ native memory: the Gemini-3-Flash router uses memory it generated itself. w/ Qwen memory: the Gemini-3-Flash router directly reuses the memory generated by the Qwen-3-30B-A3B-Instruct router, with no adaptation. The 33-agent column is the shared base pool before any expansion, so the two settings coincide there; transfer effects appear only from the first expansion ($N\geq 5$).
Setting 3 Agents 5 Agents 7 Agents 10 Agents
Gemini-3-Flash, w/ native memory 49.1% 49.7% 60.6% 62.4%
Gemini-3-Flash, w/ Qwen memory 49.1% 53.9% 60.0% 58.8%
Two observations follow. First, MonoScale remains effective on a frontier, closed-weight router: with native memory, Gemini-3-Flash improves consistently from $49.1\%$ to $62.4\%$ as the pool grows. Second, and more importantly, memory generated by the Qwen router transfers to Gemini-3-Flash without any adaptation: the transferred setting still scales from $49.1\%$ to $58.8\%$ with no collapse, and its gap to native memory stays modest and bounded (e.g., $62.4\%$ vs. $58.8\%$ at $10$ agents) rather than escalating with scale. This indicates that the memory MonoScale distills encodes transferable, router-agnostic profiles of agent capabilities and failure modes, rather than router-specific artifacts.
D.4 Scaling Beyond Ten Agents
Our main experiments scale the agent pool up to $N=10$. To probe behaviour at a larger scale, we further expand the HLE-MCQ system from $10$ to $20$ worker agents by adding ten Gemini-2.5-Pro-based workers, keeping the rest of the setup unchanged. Overall accuracy improves from $19.88\%$ at $10$ agents to $23.4\%$ at $20$ agents, indicating that at moderate scale, additional model capability and skill diversity still yield tangible gains and that the context window is not yet the dominant bottleneck. We nonetheless do not claim that the current prompt-based routing paradigm alone addresses true Web-scale settings: at hundreds or thousands of agents, context length becomes the primary bottleneck, and MonoScale would need to be combined with an agent-recommendation or hierarchical-scheduling stage that first retrieves a small relevant subset of agents before routing.
D.5 Onboarding Cost
Onboarding a new agent is a one-time cost. In our implementation, it requires roughly $50$ customized warm-up tasks with $4$ rollout trajectories each, costing about \$19.5 per agent under DeepSeek’s official API pricing. For comparison, a single full evaluation costs about \$50, and the full HLE benchmark about \$100 (roughly $5\times$ the per-agent onboarding fee). The warm-up procedure is highly parallelizable and typically completes within 10101515 minutes, so its wall-clock latency is limited. Because onboarding is performed only once per agent while the resulting agent is reused across many downstream tasks, we regard this overhead as moderate and readily amortized given the persistent gain in system capability.
D.6 Multi-Seed Variance
Full multi-seed evaluation of an LLM-based MAS is expensive, so we assess variance on a randomly sampled $40$-task GAIA subset using three independent seeds. The standard deviation of overall accuracy is $2.89$, $1.44$, and $2.89$ points at $5$, 77, and $10$ agents respectively—only about 0.50.51.21.2 tasks out of $40$. This low variance indicates that MonoScale’s gains across expansion are not an artifact of random fluctuation; accordingly, we frame the empirical trend as a stable, non-decreasing improvement rather than a strict per-point monotonicity claim.
D.7 Error Analysis of the Local Scaling Dip
The GAIA scaling curve in Figure 2 rises in overall trend but is not strictly monotonic at every point: accuracy shows a small local dip at $N=6$ ($47.3\%$ at $5$ agents versus $46.1\%$ at $6$, a difference of about $1.2$ points, or $2$ tasks out of the $165$-question GAIA validation set). Inspecting this dip, we attribute it predominantly to tool-budget exhaustion. After a new agent is integrated and warm-up completes, the increased diversity of the agent pool leads the router to decompose tasks into finer-grained subtasks and distribute them more aggressively across agents. Under a fixed per-task tool-call budget, this finer decomposition raises the number of tool calls a task requires, increasing the chance of hitting the budget limit before a conclusive answer is produced and thus causing the task to fail. These cases are concentrated in complex reasoning tasks that require multi-step sequential tool use, which are inherently most sensitive to budget constraints. We trace the root cause to the warm-up stage: its current task design does not explicitly cover budget-constrained scenarios, so the collected evidence lacks negative signals about budget exhaustion and the distilled routing memory provides little guidance on avoiding over-decomposition under a limited budget. This suggests a concrete remediation within the MonoScale framework—emphasizing budget-constrained scenarios during warm-up so that familiarization captures these failure signals and encodes them into memory. Finally, this local dip does not contradict our analysis: Theorem 4.3 guarantees a non-decrease of expected performance across expansion stages, not strict monotonicity of every finite-sample evaluation point.
D.8 Malfunctioning Agent Pool
@@
Appendix E MonoScale/Expansion-Aware Memory Schema
E.1 Overview
@@ E.1.1 Key Common Items
E.2 Planning Phase Memory
E.2.1 JSON Template
@@
E.2.2 Planning-Specific Items
@@
E.3 Coordinating Phase Memory
E.3.1 JSON Template
@@
E.3.2 Coordinating-Specific Items
@@
Appendix F Case Study Trajectories: The Anatomy of Scaling Failures and Their Cure
@@
F.1 Case 1: The Trap of Cold-Start Misjudgment (Anagram Solving)
@@
F.2 Case 2: The Trap of Brittle Workflow Links (Bird Species ID)
@@
F.3 Case 3: The MonoScale Resolution (Scientific Discovery)